OpenStudy (anonymous):

If 14 added to a certain number and the sum is multiplied by 2 the result is equal to 8 times the number decreased by 14.Find the number. If 14 added to a certain number and the sum is multiplied by 2 the result is equal to 8 times the number decreased by 14.Find the number. @Mathematics

OpenStudy (anonymous):

Let x be your number. "If 14 is added to a certain number and the sum is multiplied by 2": In algebra, this means 2(14 + x)

OpenStudy (anonymous):

How would you write, in algegbra: " 8 times the number decreased by 14" ?

OpenStudy (texaschic101):

2(14 + x) = 8x - 14 (distribute through the parenthesis) 28 + 2x = 8x - 14 2x - 8x = -14 - 28 -6x = - 42 x = -42/-6 x = 7